OpenStudy (anonymous):

please help ,Solve the following equation. (5x - 1)/2 + (-3x - 1)/4 = 9/2

OpenStudy (anonymous):

i ended up with X=10.5 / 26 as my answer but i dont think thats right...

hero (hero):

Hint: Subtract (5x-1)/2 from both sides

hero (hero):

\(\large-\frac{3x + 1}{4} = \frac{9}{2} - \frac{5x - 1}{2}\)

hero (hero):

Then combine the right side. Afterwards, cross multiply

OpenStudy (anonymous):

thanks !

hero (hero):

You should get \(-\frac{3x + 1}{4} = \frac{10 - 5x}{2}\) before cross multiplying

hero (hero):

Hopefully, you know how to properly cross multiply. Dealing with negative fractions can get tricky.

OpenStudy (goformit100):

good @unknown&confused

OpenStudy (goformit100):

you are a good learner @unknown&confused

hero (hero):

Be sure to post the value of x you get @unknown

OpenStudy (anonymous):

just a moment

hero (hero):

@unknown&confused

OpenStudy (anonymous):

x=26/41

OpenStudy (anonymous):

that correct ?

hero (hero):

Show me how you got that

OpenStudy (anonymous):